6. Let X be a uniform random variable on the interval [0, b]. Determine b so that σ = 2 √ µ.
Answer #1

For a uniform distribution, = (a + b)/2

=

When = 2 ,

2 = 4

4 = (b-a)2/12

= (b-a)2/48

(a+b)/2 = (b-a)2/48

Given that, a = 0

Therefore,

b/2 = b2/48

b = b2/24

1 = b/24

b = 24